A 219 g ball is tied to a string. It is pulled to an angle of 7.60 degrees and released to swing as a pendulum. A student with a stopwatch finds that 13 oscillations take 14.5 s .

Part A) How long is the string? Express your answer in centimeters.
Answer #1

Solution)

Time Period, T= 14.5/13 = 1.15 secs

Now, we know

L= T^2*g/ 4*3.14*3.14

L=( 1.15^2 * 9.8)/(4*3.14*3.14) = 0.328 m=32.8 cm (Ans)

========

Good luck!:)